Find the sum and product of the matrix of the given eigen values of the matrix
A=2 1 2
1 3 1
2 2 -6
1
Expert's answer
2020-12-01T06:24:59-0500

We have matrix: A=


"\begin{pmatrix}
 2 & 1 & 2 \\
 1 & 3 & 1 \\
2 & 2 & -6
\end{pmatrix}"

Find the eigenvalues of the matrix:

Associate matrix Vlasna vectors values and introduce a variable.

"\begin{pmatrix}
 2- \lambda & 1 & 2 \\
 1 & 3-\lambda & 1 \\
 2 & 2 & -6 - \lambda

\end{pmatrix}"

To do this, find the determinant of the matrix and equate this expression to zero.

"(2-\lambda)*((3 - \lambda)*(-6-\lambda)-2*1)-1*(1*(-6=\lambda)-2*2)+2*(1*1-(3-\lambda)*2)=0"

After transformations, we get:

"\lambda"3"+\lambda"2 "-31\lambda +40 = 0"

After applying the rules, it is clear that the roots of the equation are not integers. Let's calculate the approximate "\lambda" .

"\lambda"1 "=-6,60"

"\lambda"2 "=4,15"

"\lambda"3 "=1,45"

Multiply the matrix A by the matrix of its eigenvalues:

"\begin{pmatrix}
 2 & 1 & 2 \\
 1 & 3 & 1 \\
 2 & 2 & -6
\end{pmatrix}" "*" "\begin{pmatrix}
 -6,6 \\
 4,15 \\
 1,45
\end{pmatrix}" "=" "\begin{pmatrix}
 -6,15 \\
 7,3 \\
 -13,6
\end{pmatrix}"






Maybe, if I misunderstood something, I add to the report: the sum and product of the matrix A itself:

The sum of the matrix: A + A = 2A

"2\begin{pmatrix}
 2 & 1 & 2\\
 1 & 3 & 1\\
2 & 2 & -6
\end{pmatrix}=\begin{pmatrix}
 4 & 2 & 4 \\
 2 & 6 & 2\\
4 & 4 & -12 
\end{pmatrix}"


And product of the matrix: "A * A = A^2"


"A*A=\begin{pmatrix}
 2 & 1 & 2 \\
 1 & 3 & 1 \\
2 & 2 & -6
\end{pmatrix}*\begin{pmatrix}
 2 & 1 & 2 \\
 1 & 3 & 1 \\
2 & 2 & -6
\end{pmatrix}=\begin{pmatrix}
 9 & 9 & -7 \\
 7 & 12 & -1 \\
-6 & -4 & 42
\end{pmatrix}"


Multiplication is performed for each component of the matrix separately:


"a^2_11=a_11*a_11 +a_12*a_21+a_13*a_31= 2*2+1*1+2*2=9"

"a^2_12=a_11*a_12 +a_12*a_22+a_13*a_32=2*1+1*3+2*2=2+3+4=9"

"a^2_13=a_11*a_13 +a_12*a_23+a_13*a_33= 2*2+1*1+2*(-6)=-7"

"a^2_21=a_21*a_11 +a_22*a_21+a_23*a_31=1*2+3*1+1*2=7"

"a^2_22=a_21*a_12 +a_22*a_22+a_23*a_32=1*1+3*3+1*2=12"

"a^2_23=a_21*a_13 +a_22*a_23+a_23*a_33=1*2+3*1+1*(-6)=-1"

"a^2_31=a_31*a_11 +a_32*a_21+a_33*a_31= 2 *2 +2*1+(-6)*2=-6"

"a^2_32=a_31*a_12 +a_32*a_22+a_33*a_32=2*1+2*3+(-6)*2=-4"

"a^2_33=a_31*a_13 +a_32*a_23+a_33*a_33=2*2+2*1+(-6)*(-6)=42"